# Certificate-Sensitive Subset Sum: Realizing Instance Complexity

### v4

@@ Abstract
The Subset Sum problem is a classical NP-complete problem with a long-standing $O^{*}(2^{n/2})$ deterministic bound due to Horowitz and Sahni. Yet, many instances are much easier, as their complexity is governed by $U=|\Sigma(S)|$, the number of unique subset sums. We present three contributions that exploit this structure. First, we introduce the first deterministic algorithm whose runtime provably grows with the certificate size $U$, constructing all distinct subset sums in time $O(U\cdot n^{2})$, with a randomized variant achieving expected time $O(U\cdot n)$. This provides a constructive link to Instance Complexity. Second, we build a solver that combines this enumerator with a double meet-in-the-middle strategy and a Controlled Aliasing technique, achieving a worst-case runtime of $O^{*}(2^{n/2-\epsilon})$ for a constant $\epsilon>0$—the first deterministic improvement over the classical bound on all instances. Finally, we refine fine-grained hardness for Subset Sum by showing that SETH-based reductions implicitly assume collision-free instances with maximal $U$.
@@ 1 Motivation
Our work delivers a structural breakthrough on three fronts.
  • A first in certificate-sensitive complexity: This is, to our knowledge, the first deterministic algorithm for a canonical NP–complete problem whose runtime adapts to a provably minimal certificate, namely ($\Sigma(S)$). Our algorithm constructs this certificate online from scratch, rather than merely deciding feasibility.
  • Strictly faster than Horowitz–Sahni on all instances: IC-SubsetSum (formerly Beyond Worst-Case Subset Sum from [11]) incorporates a novel pruning rule that guarantees worst-case runtime $O^{*}(2^{n/2-\varepsilon})$ for a constant $\varepsilon>0$. It is the first deterministic solver to provably outperform the classical $O^{*}(2^{n/2})$ bound on every input, without exception.
  • A structural reframing of fine-grained lower bounds: Existing fine-grained reductions rely on the assumption that hard instances have $U=2^{n/2}$ (i.e., no subset sums collide). IC-SubsetSum shows that this assumption is not just technical, but structural: the runtime hardness of Subset–Sum is a direct function of its subset sum collisions.
  • A generic approach for certificate-sensitive algorithms: By reinterpreting dynamic programming as a guided certificate traversal, IC-SubsetSum offers a new design template for adaptive enumeration. This opens the door to structural algorithms for other NP–complete problems where certificate size varies widely across instances.
@@ 2 Introduction
The modern theory of algorithms is dominated by worst–case running–time guarantees. While immensely successful, the paradigm sometimes fails to explain the vast performance variance that practitioners observe between individual inputs of the same size. A complementary research line, initiated by [10], formalised Instance Complexity (IC): the intrinsic difficulty of a single instance is the bit–length of the shortest “special–case program” that decides it within a time bound $t$.
@@
Yet practitioners notice that Subset–Sum instances stemming from real data (e.g. knapsack cryptosystems, bin–packing logs) are routinely solvable in far less time. The explanation is simple: such instances exhibit enormous collision rate—many distinct subsets happen to yield identical sums. Classical analyses, parameterized solely by $n$ and $t$, overlook this structure.
@@
[…]
This work builds upon the algorithm initially introduced in the ‘Beyond Worst-Case Subset Sum’ framework [11], now formalized under the IC-SubsetSum model.
@@
  • A guaranteed worst–case improvement. We prove that IC-SubsetSum runs in $O^{*}(2^{n/2-\varepsilon})$ time for a fixed constant $\varepsilon>0$, making it strictly faster than classical meet-in-the-middle algorithms on all inputs, regardless of structure.
@@
The certificate-sensitive bound provides an exponential advantage over classical dynamic programming (DP) on instances with large numeric range but low structural complexity. For example, consider an instance $S$ with $n-1$ elements equal to 111, and one large element $L=2^{n}$. For a target $t=L$, the DP runtime is $O(n\cdot t)=O(n\cdot 2^{n})$. In contrast, the number of unique subset sums for $S$ is only $U=O(n)$, yielding a polynomial runtime of $O(U\cdot n^{2})=O(n^{3})$ for IC-SubsetSum. Furthermore, the underlying enumeration framework supports anytime and online operation, allowing for interruption and incremental updates; these features are detailed in our companion technical paper [11].
@@
This paper lays the foundational theoretical groundwork for the Certificate-Sensitive framework, building upon the algorithmic and empirical results presented in our technical companion paper [11]. The framework’s approach is shown to be broadly applicable in two other companion papers, released concurrently, that extend it to the 0–1 Knapsack problem [13, forthcoming] and the 3-SAT problem [12, forthcoming].
@@ 3 Preliminaries
Let $S=\{a_{1},\ldots,a_{n}\}\subseteq\mathbb{Z}_{>0}$ be a multiset of $n$ positive integers. We write $[n]=\{1,2,\ldots,n\}$ and let
@@
Collision entropy may be viewed as a coarse proxy for the Kolmogorov compressibility of the subset sum landscape: the more structured the set, the lower its information content.
@@
Instance Complexity [10] formalises the idea that hard problems can have easy instances—even without randomization or approximation. The IC of an instance $x$ under time bound $t$ is defined as the bit-length of the shortest program $P$ that outputs the correct answer on $x$ within time $t$. It is denoted $\mathsf{IC}_{t}(x)$.
For Subset–Sum, the list $\Sigma(S)$ plays a dual role. It is not only a certificate but also a low-complexity proxy for a correct decision procedure. Once $\Sigma(S)$ is known, any target $t$ can be resolved in time $O(1)$ via a membership query. Hence, $\Sigma(S)$ encodes a special-case program of length $O(U\log(n\cdot\max(S)))$, yielding a concrete upper bound on $\mathsf{IC}_{t}((S,t))$. When the goal is to construct a solution, the richer certificate consisting of $(\sigma,\text{bitmask})$ pairs provides a special-case program of length $O(U\cdot n)$.
In our IC discussion, we assume a time bound $t(n)=O(1)$ when $\Sigma(S)$ is stored in a hash table, or $t(n)=O(\log U)$ if stored in sorted order for binary search. Both bounds reflect the efficient nature of checking membership in a precomputed certificate set.
@@ 4.1 Unique-Subset-Sum Enumerator
Our algorithm begins by generating $\Sigma(S)$, the set of all distinct subset sums of $S$. The core challenge is to do this without duplication—i.e., without computing the same sum multiple times via different subsets. We use an ordered traversal strategy based on prefix extensions.
We emphasize that the goal of the enumerator is not merely to determine the feasibility of subset sums, but to generate the full constructive certificate as defined in Section 3. That is, for each unique subset sum $\sigma\in\Sigma(S)$, the algorithm records the lexicographically minimal bitmask of a subset realizing $\sigma$. This allows efficient reconstruction of actual witnesses and is necessary for constructive queries. Consequently, the runtime and space complexity of the enumerator will scale with the size of this richer certificate. Full implementation details and pseudocode for the enumerator are provided in Appendix A.
@@
In high-density instances, where many distinct subsets may sum to the same value, a naive ”first-prefix-wins” pruning strategy can fail. To guarantee correctness, our algorithm resolves collisions deterministically by defining a canonical prefix for each sum—this is the prefix whose index mask is lexicographically minimal among all subsets producing that sum. By retaining only this canonical representation when multiple extensions lead to the same sum, we suppress redundant work and ensure each unique sum is discovered via a single, well-defined path.
@@
The deterministic enumerator computes $\Sigma(S)$ in $O(|\Sigma(S)|\cdot n^{2})$ worst-case time and requires $O(|\Sigma(S)|\cdot n)$ space.
@@
The runtime is dominated by the prefix extension and deduplication process. There are $U=|\Sigma(S)|$ unique sums, or ”states,” to be discovered. From each state, the algorithm attempts to extend its prefix with one of at most $n$ elements, resulting in $O(U\cdot n)$ extension attempts. In the event of a sum collision, the algorithm must select a canonical prefix by comparing two $n$-bitmasks, which takes $O(n)$ time. The total runtime is therefore $O(U\cdot n)\times O(n)=O(U\cdot n^{2})$. The space complexity is determined by storing the $n$-bitmask for each of the $U$ states, resulting in $O(U\cdot n)$ space usage. ∎
@@
A randomized variant of the enumerator computes $\Sigma(S)$ in expected time $O(|\Sigma(S)|\cdot n)$.
@@
The randomized variant replaces the $O(n)$ prefix comparison with an expected $O(1)$ comparison using polynomial hashing (e.g., from a 2-universal hashing family) with a randomly chosen seed. This ensures that, with high probability, no two distinct bitmasks collide. Thus, the expected comparison cost per extension remains $O(1)$ even under adversarial inputs. ∎
Remark 1.
The expected runtime bound for the randomized variant assumes a standard cryptographic model in which the hash function is drawn from a universal family. This ensures robustness against adversarially chosen inputs by keeping the probability of bitmask collisions negligible. To ensure a negligible collision probability (e.g., $1/\mathrm{poly}(n,U)$), the polynomial hash function is evaluated over a prime field of size $\mathrm{poly}(n,U)$.

### v6

@@ Abstract
The Subset Sum problem is a classical NP-complete problem with a long-standing $O^{*}(2^{n/2})$ deterministic bound due to Horowitz and Sahni. We present results at two distinct levels of generality.
First (instance-sensitive bound), we introduce, to our knowledge, the first deterministic algorithm whose runtime provably scales with the certificate size $U=|\Sigma(S)|$, the number of distinct subset sums. Our enumerator constructs all such sums in time $O(U\cdot n^{2})$, with a randomized variant achieving expected time $O(U\cdot n)$. This provides a constructive link to Instance Complexity by tying runtime to the size of an information-theoretically minimal certificate.
Second (unconditional worst-case bound), by combining this enumerator with a double meet-in-the-middle strategy and a Controlled Aliasing technique that enforces a simple canonical-normal-form (CNF) expansion policy on aliased states, we obtain a deterministic solver running in $O^{*}(2^{n/2-\varepsilon})$ time with $\varepsilon=\log_{2}(\frac{4}{3})$—the first unconditional deterministic improvement over the classical $O^{*}(2^{n/2})$ bound for all sufficiently large $n$.
Finally, we refine fine-grained hardness for Subset Sum by making explicit the structural regime (high collision entropy / near collision-free) implicitly assumed by SETH-based reductions, i.e., instances with near-maximal $U$.
@@ 1 Motivation
This work offers four primary contributions, three of which are structural in nature.
  • Instance-sensitive bound via certificate size $U$: To our knowledge, this is the first deterministic algorithm for a canonical NP–complete problem whose runtime provably adapts to a minimal constructive certificate, namely $\Sigma(S)$. The algorithm constructs this certificate online from scratch, in deterministic time $O(U\cdot n^{2})$ (expected $O(U\cdot n)$ randomized), rather than merely deciding feasibility.
  • Unconditional worst-case improvement: By combining the enumerator with a double meet-in-the-middle strategy and Controlled Aliasing under Canonical Normal Form (CNF)—i.e., a simple count-based expansion policy at the alias indices (§5.2, App. E)—IC-SubsetSum achieves a worst-case runtime $O^{*}(2^{\,n/2-\varepsilon})$ with $\varepsilon=\log_{2}\!\left(\tfrac{4}{3}\right)$ for all inputs, the first deterministic improvement over the classical $O^{*}(2^{n/2})$ Horowitz–Sahni bound in nearly 50 years.
  • Structural reframing of fine-grained hardness: Many fine-grained reductions to Subset–Sum implicitly target high-entropy, near collision-free instances (i.e., $U\approx 2^{n}$). IC-SubsetSum makes this structural dependency explicit, showing that hardness in the standard reductions aligns with the regime of near-maximal distinct sums.
  • A generic design template for certificate-sensitive algorithms: By reinterpreting dynamic programming as guided certificate traversal, IC-SubsetSum offers a broadly applicable approach to adaptive enumeration, potentially extending to other NP–complete problems with wide certificate-size variance.
@@ 2 Introduction
The modern theory of algorithms is dominated by worst–case running–time guarantees. While immensely successful, the paradigm sometimes fails to explain the vast performance variance that practitioners observe between individual inputs of the same size. A complementary research line, initiated by Orponen–Ko–Schöning–Watanabe [10], formalised Instance Complexity (IC): the intrinsic difficulty of a single instance is the bit–length of the shortest “special–case program“ that decides it within a time bound $t$.
@@
Yet practitioners observe that real-data instances (e.g., knapsack crypto, bin-packing logs) are often much easier: many different subsets collide to the same value, so the number of distinct subset sums $U=|\Sigma(S)|$ is far smaller than $2^{n}$. This additive redundancy—small $U$ due to duplicates, near–arithmetic progressions, or other structure—is exactly what analyses that track only $n$ and $t$ overlook.
@@
We address this gap. Let $\Sigma(S)=\{\sum_{i\in T}a_{i}:T\subseteq[n]\}$ be the set of distinct subset sums and write $U=|\Sigma(S)|$. Because deciding $(S,t)$ reduces to a simple membership query once $\Sigma(S)$ is known, this list is an information–theoretically minimal certificate. Our new algorithm IC-SubsetSum deterministically enumerates every element of $\Sigma(S)$ exactly once, prunes duplicates on the fly, and halts in deterministic worst-case time $O(U\cdot n^{2})$. We further show this can be improved to an expected time of $O(U\cdot n)$ using randomization. Thus its runtime is provably bounded by a function of the certificate size, marking the first algorithm for Subset–Sum whose complexity scales with the true structural difficulty of the input. We build upon the algorithm initially introduced in the ‘Beyond Worst-Case Subset Sum’ framework [13], now formalized under the IC-SubsetSum model.
@@
  • A guaranteed worst–case improvement. We prove that IC-SubsetSum runs in $O^{*}(2^{\,n/2-\varepsilon})$ time, with $\varepsilon=\log_{2}(\frac{4}{3})$, by combining double meet-in-the-middle with Controlled Aliasing under a fixed count-based canonical expansion policy (CNF; see §5.2, App. E). This yields a deterministic speedup over classical meet-in-the-middle on all inputs.
@@
The certificate-sensitive bound provides an exponential advantage over classical dynamic programming (DP) on instances with large numeric range but low structural complexity. For example, consider an instance $S$ with n1n-1 elements equal to $1$, and one large element $L=2^{n}$. For a target $t=L$, the DP runtime is $O(n\cdot t)=O(n\cdot 2^{n})$. In contrast, the number of unique subset sums for $S$ is only $U=O(n)$, yielding a polynomial runtime of $O(U\cdot n^{2})=O(n^{3})$ for IC-SubsetSum. Furthermore, the underlying enumeration framework supports anytime and online operation, allowing for interruption and incremental updates; these features are detailed in our companion technical paper [13].
Unlike prior algorithms whose performance depends on less precise proxies (e.g., numeric range, pseudo-polynomial bounds, or number of solutions), our solver’s runtime is provably governed exactly by $U=|\Sigma(S)|$ for every instance—the first such result treating $U$ as a formal instance-complexity certificate.
@@
This paper lays the foundational theoretical groundwork for the Certificate-Sensitive framework, building upon the algorithmic and empirical results presented in our technical companion paper [13]. The framework’s approach is shown to be broadly applicable in two other companion papers, released concurrently, that extend it to the 0–1 Knapsack problem [11, forthcoming] and the 3-SAT problem [12, forthcoming].
@@ 3 Preliminaries
Computational Model.
We adopt the standard Word–RAM model with word size $w=\Theta(\log M)$, where $M:=\max\{\sum_{a\in S}a,\;t\}$, so that arithmetic, comparisons, and memory accesses on $w$-bit words take $O(1)$ time. Equivalently, $w$ is large enough that every subset sum and $t$ fit in a single word, so additions and comparisons on sums are unit-cost. All logarithms are base two unless stated otherwise.
Randomized Model.
Our randomized variants rely on standard 2-universal hashing. We assume that evaluating a hash function and handling collisions can be done in expected $O(1)$ time in the Word–RAM model. For any $x\neq y$, $\Pr_{h\sim\mathcal{H}}[h(x)=h(y)]\leq 1/n^{c}$ for a fixed constant $c\geq 1$, and expectations are taken over the random draw of $h$. Collisions are resolved by exact bitmask comparison, ensuring Las Vegas correctness: the output is always correct, and the runtime bound holds in expectation.
Asymptotic Notation.
We use $O^{*}(\cdot)$ and $o^{*}(\cdot)$ to suppress factors polynomial in $n$ and $\log M$, unless explicitly shown.
@@
Let $(S,t)$ be a Subset Sum instance, where $S=\{a_{1},\ldots,a_{n}\}\subseteq\mathbb{Z}_{\geq 0}$ is a multiset of $n$ nonnegative integers and $t\in\mathbb{Z}_{\geq 0}$ is a nonnegative target. We write $n]=\{1,2,\ldots,n\}$ and let
@@
Collision entropy may be informally viewed as a coarse proxy for the Kolmogorov compressibility of the subset sum landscape: the more structured the set, the lower its information content.
Note. We use the unnormalized form of collision entropy $H_{c}(S)=n-\log_{2}U$, which is standard in information theory; the normalized form $H_{c}(S)/n$ differs only by a multiplicative constant and is not required for our purposes.
@@
Given a target $t$, if a solution exists, a witness is returned in additional $O(n)$ time from our maintained encodings. Unless otherwise stated (e.g., in decision-variant optimizations), all runtime bounds are for full constructive-certificate enumeration of $\Sigma(S)$, in which each $\sigma\in\Sigma(S)$ is stored with its canonical (lexicographically minimal) witness bitmask. This invariant is maintained throughout, ensuring correctness even on high-density instances.
@@
Orponen–Ko–Schöning–Watanabe Instance Complexity [10] formalises the idea that hard problems can have easy instances—even without randomization or approximation. The IC of an instance $x$ under time bound $t$ is defined as the bit-length of the shortest program $P$ that outputs the correct answer on $x$ within time $t$. It is denoted $\mathsf{IC}_{t}(x)$.
For Subset–Sum, the list $\Sigma(S)$ plays a dual role. It is not only a certificate but also a low-complexity proxy for a correct decision procedure. Once $\Sigma(S)$ is known, any target $t$ can be resolved in time $O(1)$ via a membership query. Hence, $\Sigma(S)$ encodes a special-case program of length $O(U\log(n\cdot\max(S)))$, yielding a concrete upper bound on $\mathsf{IC}_{t}((S,t))$. When the goal is to construct a solution, the richer certificate consisting of $(\sigma,\text{bitmask})$ pairs provides a special-case program of length $O(U\cdot n)$. For a formal IC program that satisfies partial correctness on all inputs, we pair $\Sigma(S)$ with an input-checking wrapper; see §4.3.
Operational Parameters for Halves.
When $S$ is split into two halves $\ell_{0},\ell_{1}$ (sizes within $\pm 1$), we write $U_{0}:=|\Sigma(\ell_{0})|$, $U_{1}:=|\Sigma(\ell_{1})|$, and $\widehat{U}:=\max\{U_{0},U_{1}\}$. When we state bounds that scale with $U$, we refer to full enumeration of $\Sigma(S)$; for meet-in-the-middle solvers operating on halves, the governing parameter is $\widehat{U}$.
Controlled Aliasing Convention.
In the Controlled Aliasing rule, each half designates an arbitrary ordered pair $(x_{0},x_{1})$ of distinct elements at indices $(i_{0},i_{1})$ and treats occurrences of $x_{1}$ as $x_{0}$ during subset-sum generation for that half. We will key memo entries by a 2-lane identifier $(\tilde{\sigma},\chi)$, where $\tilde{\sigma}$ is the aliased sum and $\chi\in\{0,1\}$ records whether $x_{1}$ is present; witness selection remains lexicographic within each lane. Unless otherwise stated, the choice of alias pairs is independent of the target $t$ (and of any randomness), and all guarantees hold for all targets $t$.
Count-based canonical normal form (CNF). We restrict expansions to canonical aliased states using a fixed, target-independent policy: for a bitmask $b$, let $c(b):=b[i_{0}]+b[i_{1}]\in\{0,1,2\}$ and $\chi(b):=b[i_{1}]\in\{0,1\}$. A newly discovered state with mask $b$ is enqueued for extension iff
[…]
This CNF policy preserves completeness and enables the $3/4$ expansion accounting; see §5.2 and App. E (Lemmas 910).
Trivial and Degenerate Cases.
Instances with $n=0$ or $t=0$ are handled in $O(1)$ time. If $U=1$ (e.g., all subset sums collide), enumeration and decision terminate immediately. Duplicates in $S$ are handled natively by our canonical memoization and require no additional promises.
Instance Assumptions.
We consider the Subset Sum problem over a multiset $S=\{a_{1},\dots,a_{n}\}$ of nonnegative integers ($a_{i}\in\mathbb{Z}_{\geq 0}$) and an integer target $t\geq 0$. Elements are not required to be distinct; duplicates are treated as separate items. We assume without loss of generality that $n\geq 1$ and that $t\leq\sum_{a\in S}a$ (instances with $t<0$ or $t>\sum S$ are decided trivially). Zero-valued elements may appear and are processed without special handling. While they do not generate new sum values (leaving $U=|\Sigma(S)|$ unchanged), they can introduce additional witness subsets for existing sums. The algorithm’s correctness is unaffected, as its canonicalization logic correctly identifies the lexicographically minimal witness in all cases. Even if $t=0$, our algorithms enumerate the full constructive certificate for $\Sigma(S)$, including the canonical witness for every achievable sum, unless the run is terminated early. All algorithms operate under the Word–RAM model described above, with $w=\Theta(\log M)$ where $M:=\max\{\sum_{a\in S}a,\;t\}$.
@@ 4.1 Unique-Subset-Sum Enumerator
Our algorithm begins by generating $\Sigma(S)$, the set of all distinct subset sums of $S$. The core challenge is to do this without duplication—i.e., without computing the same sum multiple times via different subsets. We use an ordered traversal strategy based on prefix extensions together with the following invariant.
Invariant 1.
(Suppression-on-overwrite). Whenever a canonical representative for a numeric sum $\sigma$ is replaced by a lexicographically smaller mask, the displaced representative is marked non-expandable and never extended further. This guarantees each canonical sum is expanded exactly once.
A minimal recursive formulation (for clarity).
Before the optimized column-wise implementation (Appendix A), it helps to view the algorithm abstractly:
Algorithm 1 AbstractEnumerateUnique(S)(S)
1:Memo { $\leftarrow\{\,0\mapsto\texttt{empty\_bitmask}\,\}$
2:function Extend($\sigma,b,j$) $\triangleright$ $b$ is a bitmask; $j$ is next index
3:  for $i=j$ to $n$ do
4:   $\sigma^{\prime}\leftarrow\sigma+a_{i}$; $b^{\prime}\leftarrow b$ with bit $i$ set
5:   if $\sigma^{\prime}\notin\texttt{Memo}$ then
6:     Memo[$\sigma^{\prime}$] $\leftarrow b^{\prime}$
7:     Extend($\sigma^{\prime},b^{\prime},i{+}1$)
8:   else if IsLexicographicallySmaller($b^{\prime}$, Memo[$\sigma^{\prime}$]) then
9:     Memo[$\sigma^{\prime}$] $\leftarrow b^{\prime}$ $\triangleright$ overwrite with lex-min representative
10:     Extend($\sigma^{\prime},b^{\prime},i{+}1$)
11:Extend($0,\texttt{empty\_bitmask},1$)
12:return Memo
This recursive view exposes the two core operations: (i) emit-or-overwrite a representative for a sum, and (ii) extend only from the (current) representative. The optimized frontier-based implementation in Appendix A enforces Invariant 1 explicitly via a doNotExtend flag so that an overwritten representative is never expanded again.
We emphasize that the goal of the enumerator is not merely to determine feasibility, but to generate the full constructive certificate as defined in Section 3. For each unique subset sum $\sigma\in\Sigma(S)$, the algorithm records the lexicographically minimal bitmask of a subset realizing $\sigma$. This enables efficient reconstruction of witnesses and is necessary for correctness in constructive queries. Consequently, the runtime and space of the enumerator scale with the size of this richer certificate. Full implementation details and pseudocode are provided in Appendix A. (When Controlled Aliasing is enabled, the same enumeration principles apply per half with 2-lane memoization; see §5.2 and App. E.)
@@
In high-density instances, where many distinct subsets may sum to the same value, a naive “first-prefix-wins” pruning strategy can fail. To guarantee correctness, our algorithm resolves collisions deterministically by defining a canonical prefix for each sum—the prefix whose index mask is lexicographically minimal among all subsets producing that sum (compare bits from $1$ to $n$, preferring $0<1$ at the first difference). By retaining only this canonical representation when multiple extensions lead to the same sum, we suppress redundant work and ensure each unique sum is discovered via a single, well-defined path.
@@
Separation of concerns.
Conceptually, the enumerator factors into three orthogonal pieces: (i) enumeration (systematically proposing extensions), (ii) canonicality tests (lexicographic comparisons to select representatives), and (iii) memoization (storing one representative per numeric sum). Invariant 1 links (ii) and (i), ensuring that only the current representative ever generates successors.
@@
Under the computational model of Section 3, the deterministic enumerator computes all $U=|\Sigma(S)|$ unique subset sums in $O(U\cdot n^{2})$ time and $O(U\cdot n)$ space.
@@
The runtime is dominated by the prefix extension and deduplication process. From each of the $U$ states, the algorithm attempts to extend its prefix with at most $n$ elements, resulting in $O(U\cdot n)$ extension attempts. In the event of a sum collision, the algorithm must compare the existing and candidate $n$-bitmasks to select the canonical representative, costing $O(n)$ time. Invariant 1 ensures that an overwritten representative is never expanded further. Multiplying $O(U\cdot n)$ extension attempts by $O(n)$ comparison cost gives the stated $O(U\cdot n^{2})$ runtime bound; space is $O(U\cdot n)$ for storing all canonical bitmasks. ∎
Remark 1 (Degenerate Cases).
The $O(U\cdot n^{2})$ bound in Theorem 1 is a worst-case guarantee. For instances with very low structural complexity (e.g., $U=1$ for an empty set, or $U=O(n)$ for a set of identical elements), this bound correctly implies a fast, polynomial runtime.
@@
Let $U=|\Sigma(S)|$. Assume access to a $2$-universal hash family $\mathcal{H}$ with $O(1)$ evaluation time. Then a randomized variant of the enumerator computes all $U$ unique subset sums in expected time $O(U\cdot n)$, where the expectation is over the random draw of $h\sim\mathcal{H}$ as specified in Section 3.
@@
Assume the randomized model of Section 3. With $h$ drawn uniformly from a $2$-universal family, the probability that two distinct bitmasks yield the same hash is at most $1/\mathrm{poly}(n,U)$. Consequently, the full $O(n)$ lexicographic comparison is needed only on (rare) hash-collision events, giving an expected $O(1)$ time per canonicality check (Las Vegas correctness). Since there are $O(U\cdot n)$ extension attempts (Theorem 1), this yields the stated expected runtime bound. ∎